# Impartial Games: a Challenge for Reinforcement Learning

### v5

@@ 5.2 Experiment setup and results
This section details the experimental configurations and presents the results derived from training our AlphaZero-style algorithm on various Nim games. The performance of the trained agents is subsequently analyzed in depth, drawing upon the ’champion’ and ’expert’ perspectives introduced earlier.
@@
The specific variant of Nim studied is described in Section 2, where each heap initially contains an odd number of counters. For instance, the initial board for 5-heap Nim (as depicted conceptually in Fig. 1) consists of heap sizes [1,3,5,7,9][1,3,5,7,9][ 1 , 3 , 5 , 7 , 9 ]. The board representation provided as input to the neural network encodes the counters using unitary numbers (1 for present, 0 for removed), with -1 acting as a delimiter between distinct heaps. This provides a variable-length bitstring representation to the LSTM.
The state and action spaces for the Nim variants are as follows:
@@
On average, the game length when moves are selected randomly by two players is approximately 10 for 5-heap Nim, 13 for 6-heap Nim, and 16 for 7-heap Nim. During training, we collected 100 episodes of self-play interaction data in the form of $(s,\pi,r)$ tuples (state, MCTS-derived policy target, game outcome reward) for updating the neural networks. The policy network was trained using cross-entropy loss, and the value network using mean squared error (MSE). The detailed performance analysis under the champion and expert perspectives, leveraging this experimental setup, is presented in the following sections.
@@ 5.2.1 Champion measure by Elo rating
The Elo rating system, widely adopted in competitive two-player games such as chess, provides a robust methodology for ranking players based on their relative competitive strength. This constitutes a ’champion measure’, reflecting an entity’s ability to consistently defeat opponents. AlphaZero itself utilized Elo ratings to quantify its performance against formidable counterparts like AlphaGo Zero, AlphaGo Lee, and Stockfish. Given the absence of a pre-established competitive Nim agent with an associated Elo rating, we implemented a variation of the self-play Elo rating system [38]. This approach enables us to continuously monitor the relative strength of our agent and track its training progress, with the rating dynamically updated against all its historical predecessors to ensure an accurate reflection of its evolving competitiveness.
In our self-play Elo rating system, each newly trained agent is initialized with a baseline score of 1000. Following each training iteration, the newly trained agent is integrated into a repository of previously trained agents. Its Elo rating is then computed by conducting evaluation matches against all agents currently stored in this system. For any given match, let the agent undergoing training be denoted as Player A, and its opponent, one of its predecessors, as Player B. The expected score for Player A ($E_{A}$), representing its probability of winning the match, is calculated using the logistic function:
@@ 5.2.2 Expert measure by the accuracy of PV-NN
Beyond assessing an agent’s ’champion’ capabilities through Elo ratings, a deeper understanding of its ’expert’ mastery necessitates evaluating the intrinsic accuracy of its PV-NN. From an expert perspective, the policy network’s action probability distribution should accurately reflect optimal moves. In Nim, the mathematically provable optimal moves are uniquely determined by the nim-sum property. Consequently, the policy network’s accuracy is evaluated by comparing the probability assigned to the optimal move (as given by the nim-sum) against its actual prediction, specifically focusing on the most probable move. A similar policy evaluation metric, termed "Policy Top 1 Accuracy," has been employed in related works [14].
@@
The policy network learns from the search probabilities $\bm{\pi}(\textbf{a}|s)$ derived from the MCTS, as detailed in Formula 7. The value network learns from the actual game outcomes recorded during self-play. To further analyze the learning dynamics of these components, we tracked their respective training losses throughout the process, presented in Figure 10. It is evident that both policy and value networks can progressively fit their respective targets, which contributes to the agent’s increasing competitiveness (as reflected by Elo). However, this convergence in loss does not translate into expert-level accuracy. The inherent difficulty for the policy network to accurately learn the complex heuristics and for the value network to precisely model expected outcomes intensifies with increasing board size. This foundational problem significantly impedes the agent’s progression towards true expert-level play.
@@
The consistently decreasing training losses of the PV-NN, coupled with the rising Elo ratings of the agents, strongly suggest that the agent is learning to become a ’champion’ within its self-play ecosystem. However, the concurrent and significant decline in the PV-NN’s accuracy as the board size increases underscores a fundamental challenge in achieving ’expert’ mastery. On 7-heap Nim, the PV-NN appears capable of memorizing the heuristics and game outcomes from frequently encountered self-play data. Crucially, this rote memorization is insufficient to provide the robust, generalizable guidance needed by the MCTS to explore the vast state space effectively and form a truly constructive policy improvement loop. This highlights a critical disconnect where fitting training data does not equate to learning the underlying mathematical principles of the game.
5.2.3 Analysis on some Nim positions
While the preceding sections provided a high-level performance overview across various Nim board sizes, this section delves into specific case studies to provide a fine-grained analysis of the algorithm’s behavior. We evaluate the performance of our trained AlphaZero agent on selected initial and intermediate positions for 5, 6, and 7 heaps Nim. This approach aligns with methodologies used in analyzing game AI, such as the statistical analysis of LCZero’s performance on particular Chess positions in Section 3.
The analysis of the initial board position for a 5-heap Nim game, as evaluated by our trained PV-NN, is presented in Figure 11. All values are computed from the perspective of the current player. Moves are denoted by a heap label (e.g., ’e’) followed by the number of counters removed from that heap (e.g., ’e9’ signifies removing 9 counters from heap ’e’). In this specific initial position, the policy network assigns an exceptionally high prior probability of 97.9% to the winning move (e9), and the value network accurately estimates the resulting state’s advantage for the current player. While this strong, accurate prioritization of the optimal move is highly desirable for efficient search, it also underscores a potential vulnerability: an ungrounded overconfidence in an erroneous prior could lead to catastrophic and irrecoverable search misdirections, where MCTS simulations are unable to correct the initial bias.
@@ 5.2.3 Analysis on some Nim positions
In contrast, Figure 12 illustrates an intermediate 5-heap Nim position where the MCTS is misled. The policy network exhibits an extreme overconfidence in the losing move ’e8’, assigning it an overwhelming prior probability of 97.4%. Crucially, the true winning move for this position is ’e7’, which is not among the top-prioritized moves by the policy network. Despite the value network correctly predicting that taking ’e8’ leads to a disadvantageous position for the current player (V-value of -0.99), the algorithm entirely fails to identify the winning move ’e7’ even after an extensive 4 million MCTS simulations. The posterior probability of ’e7’ actually decreases with more simulations (Table LABEL:tab:heaps5_2), indicating the MCTS is reinforcing the incorrect prior rather than correcting it. This failure is even more pronounced than the challenges observed in certain Tesuji problems in Go, where millions of simulations might eventually yield the best action [32].
@@
Turning to the 6-heap Nim, Figure 13 shows the analysis for its initial position. Here, the policy network strongly favors the winning move ’b2’, and the value network confidently and accurately evaluates the resulting positions. This indicates that for the initial, frequently encountered states of moderately sized boards, the PV-NN can still provide effective guidance.
@@
Figure 14: In this position, the only winning move is ’f10’. The policy network’s top 4 prior probabilities are assigned to losing moves, initially hindering the search. While the value network sometimes mispredicts the outcome of certain losing moves (e.g., ’f9’ having a high empirical win probability), the MCTS eventually identifies and strongly favors the winning move ’f10’ after 65,536 simulations, and its confidence is further boosted with more simulations.
@@
Figure 15: For this 7-heap Nim position, the policy network’s prior probabilities are nearly uniform, with the top 4 assigned to losing moves. The value network’s evaluations are near zero (approximating 50% win probability). Despite extensive MCTS, the true winning moves (’e7’, ’f7’, ’g11’) are not recognized, and their posterior probabilities remain negligibly low.
@@
Across Nim variants with 5, 6, and 7 heaps, our analysis reveals a duality: positions where the PV-NN succeeds in evaluation alongside positions where it fundamentally fails. The high accuracy of predictions on initial positions of smaller boards suggests a greater learned knowledge for frequently encountered states. However, the confidence in prioritizing optimal moves deteriorates, and the ability to accurately evaluate general positions escalates in difficulty as the board size and associated state space grow.
@@
On relatively small boards, the underlying parity computations for nim-sum remain computationally feasible, and the neural networks can approximate this function to a reasonable degree. However, mastering Nim on larger boards becomes dramatically more challenging because identifying the optimal move and accurately evaluating large board positions critically entails solving complex parity-related issues using neural networks. The immense state space necessitates the PV-NNs to generalize effectively to unseen states, which, as previously argued, proves exceptionally difficult for functions with high complexity like parity. Furthermore, the inherent difficulties are compounded by the unknown and unpredictable data distribution generated by self-play reinforcement learning, as the characteristics of training data profoundly influence a neural network’s capacity to model perfect parity functions [13, 12]. The persistent problem of noisy labels, a known challenge in self-play RL, further exacerbates the difficulty for the PV-NN to learn an accurate parity function [43]. Collectively, these complexities represent a fundamental and significant challenge to contemporary reinforcement learning algorithms in their pursuit of mastering Nim. Our results empirically demonstrate that in a real-world self-play RL setting, the parity function presents a more substantial obstacle for an RL agent to achieve expert-level proficiency in Nim. Given that all impartial games can be theoretically translated into an equivalent Nim game, these identified challenges naturally extend to the broader class of impartial games.

### v6

@@ 5.2 Experiment setup and results
This section details the experimental configurations and presents the results derived from training our AlphaZero-style algorithm on various Nim games. The performance of the trained agents is subsequently analysed in depth, drawing upon the ’champion’ and ’expert’ perspectives introduced earlier.
@@
The specific variant of Nim studied is described in Section 2, where each heap initially contains an odd number of counters. For instance, the initial board for 5-heap Nim (as depicted conceptually in Fig. 1) consists of heap sizes $[1,3,5,7,9]$. The board representation provided as input to the neural network encodes the counters using unitary numbers (1 for present, 0 for removed), with -1 acting as a delimiter between distinct heaps. This provides a variable-length bitstring representation to the LSTM.
As shown in Table 3, the state and action space sizes grow rapidly with the number of heaps. On average, the game length when moves are selected randomly by two players is approximately 10 for 5-heap Nim, 13 for 6-heap Nim, and 16 for 7-heap Nim. During training, we collected 100 episodes of self-play interaction data in the form of $(s,\pi,r)$ tuples (state, MCTS-derived policy target, game outcome reward) for updating the neural networks. The policy network was trained using cross-entropy loss, and the value network using mean squared error (MSE). The detailed performance analysis under the champion and expert perspectives, leveraging this experimental setup, is presented in the following sections.
@@ 5.2.1 Champion measure by Elo rating
The Elo rating system, widely adopted in competitive two-player games such as chess, provides a robust methodology for ranking players based on their relative competitive strength. This constitutes a ’champion measure’, reflecting an entity’s ability to consistently defeat opponents. AlphaZero itself utilised Elo ratings to quantify its performance against formidable counterparts like AlphaGo Zero, AlphaGo Lee, and Stockfish. Given the absence of a pre-established competitive Nim agent with an associated Elo rating, we implemented a variation of the self-play Elo rating system (Tian et al., 2019). This approach enables us to continuously monitor the relative strength of our agent and track its training progress, with the rating dynamically updated against all its historical predecessors to ensure an accurate reflection of its evolving competitiveness.
In our self-play Elo rating system, each newly trained agent is initialised with a baseline score of 1000. Following each training iteration, the newly trained agent is integrated into a repository of previously trained agents. Its Elo rating is then computed by conducting evaluation matches against all agents currently stored in this system. For any given match, let the agent undergoing training be denoted as Player A, and its opponent, one of its predecessors, as Player B. The expected score for Player A ($E_{A}$), representing its probability of winning the match, is calculated using the logistic function:
@@ 5.2.2 Expert measure by the accuracy of PV-NN
Beyond assessing an agent’s ’champion’ capabilities through Elo ratings, a deeper understanding of its ’expert’ mastery necessitates evaluating the intrinsic accuracy of its PV-NN. From an expert perspective, the policy network’s action probability distribution should accurately reflect optimal moves. In Nim, the mathematically provable optimal moves are uniquely determined by the nim-sum property. Consequently, the policy network’s accuracy is evaluated by comparing the probability assigned to the optimal move (as given by the nim-sum) against its actual prediction, specifically focusing on the most probable move. A similar policy evaluation metric, termed "Policy Top 1 Accuracy," has been employed in related works (Danihelka et al., 2022).
@@
The policy network learns from the search probabilities $\bm{\pi}(\textbf{a}|s)$ derived from the MCTS, as detailed in Formula 8. The value network learns from the actual game outcomes recorded during self-play. To further analyse the learning dynamics of these components, we tracked their respective training losses throughout the process, presented in Figure 10. It is evident that both policy and value networks can progressively fit their respective targets, which contributes to the agent’s increasing competitiveness (as reflected by Elo). However, this convergence in loss does not translate into expert-level accuracy. The inherent difficulty for the policy network to accurately learn the complex heuristics and for the value network to precisely model expected outcomes intensifies with increasing board size. This foundational problem significantly impedes the agent’s progression towards true expert-level play.
@@
The consistently decreasing training losses of the PV-NN, coupled with the rising Elo ratings of the agents, strongly suggest that the agent is learning to become a ’champion’ within its self-play ecosystem. However, the concurrent and significant decline in the PV-NN’s accuracy as the board size increases underscores a fundamental challenge in achieving ’expert’ mastery. On 7-heap Nim, the PV-NN appears capable of memorising the heuristics and game outcomes from frequently encountered self-play data. Crucially, this rote memorisation is insufficient to provide the robust, generalizable guidance needed by the MCTS to explore the vast state space effectively and form a truly constructive policy improvement loop. This highlights a critical disconnect where fitting training data does not equate to learning the underlying mathematical principles of the game.
5.2.3 Analysis on representative Nim positions
While the preceding sections provided a high-level performance overview across various Nim board sizes, this section delves into specific case studies to provide a fine-grained analysis of the algorithm’s behaviour. We evaluate the performance of our trained AlphaZero agent on selected initial and intermediate positions for 5, 6, and 7-heap Nim. This approach aligns with methodologies used in analysing game AI, such as the statistical analysis of LCZero’s performance on particular Chess positions in Section 3.
The analysis of the initial board position for a 5-heap Nim game, as evaluated by our trained PV-NN, is presented in Figure 11. All values are computed from the perspective of the current player. Moves are denoted by a heap label (e.g., ’e’) followed by the number of counters removed from that heap (e.g., ’e9’ signifies removing 9 counters from heap ’e’). In this specific initial position, the policy network assigns an exceptionally high prior probability of 97.9% to the winning move (e9), and the value network accurately estimates the resulting state’s advantage for the current player. While this strong, accurate prioritisation of the optimal move is highly desirable for efficient search, it also underscores a potential vulnerability: an ungrounded overconfidence in an erroneous prior could lead to catastrophic and irrecoverable search misdirections, where MCTS simulations are unable to correct the initial bias.
@@ 5.2.3 Analysis on representative Nim positions
In contrast, Figure 12 illustrates an intermediate 5-heap Nim position where the MCTS is misled. The policy network exhibits an extreme overconfidence in the losing move ’e8’, assigning it an overwhelming prior probability of 97.4%. Crucially, the true winning move for this position is ’e7’, which is not among the top-prioritised moves by the policy network. Despite the value network correctly predicting that taking ’e8’ leads to a disadvantageous position for the current player (V-value of -0.99), the algorithm entirely fails to identify the winning move ’e7’ even after an extensive 4 million MCTS simulations. The posterior probability of ’e7’ actually decreases with more simulations (Table LABEL:tab:heaps5_2), indicating the MCTS is reinforcing the incorrect prior rather than correcting it. This failure is even more pronounced than the challenges observed in certain Tesuji problems in Go, where millions of simulations might eventually yield the best action (Shih et al., 2022).
@@
Turning to the 6-heap Nim, Figure 13 shows the analysis for its initial position. Here, the policy network strongly favours the winning move ’b2’, and the value network confidently and accurately evaluates the resulting positions. This indicates that for the initial, frequently encountered states of moderately sized boards, the PV-NN can still provide effective guidance.
@@
Figure 14: In this position, the only winning move is ’f10’. The policy network’s top 4 prior probabilities are assigned to losing moves, initially hindering the search. While the value network sometimes mispredicts the outcome of certain losing moves (e.g., ’f9’ having a high empirical win probability), the MCTS eventually identifies and strongly favours the winning move ’f10’ after 65,536 simulations, and its confidence is further boosted with more simulations.
@@
Figure 15: For this 7-heap Nim position, the policy network’s prior probabilities are nearly uniform, with the top 4 assigned to losing moves. The value network’s evaluations are near zero (approximating 50% win probability). Despite extensive MCTS, the true winning moves (’e7’, ’f7’, ’g11’) are not recognised, and their posterior probabilities remain negligibly low.
@@
Across Nim variants with 5, 6, and 7 heaps, our analysis reveals a duality: positions where the PV-NN succeeds in evaluation alongside positions where it fundamentally fails. The high accuracy of predictions on initial positions of smaller boards suggests a greater learned knowledge for frequently encountered states. However, the confidence in prioritising optimal moves deteriorates, and the ability to accurately evaluate general positions escalates in difficulty as the board size and associated state space grow.
@@
On relatively small boards, the underlying parity computations for nim-sum remain computationally feasible, and the neural networks can approximate this function to a reasonable degree. However, mastering Nim on larger boards becomes dramatically more challenging because identifying the optimal move and accurately evaluating large board positions critically entails solving complex parity-related issues using neural networks. The immense state space necessitates the PV-NNs to generalise effectively to unseen states, which, as previously argued, proves exceptionally difficult for functions with high complexity like parity. Furthermore, the inherent difficulties are compounded by the unknown and unpredictable data distribution generated by self-play reinforcement learning, as the characteristics of training data profoundly influence a neural network’s capacity to model perfect parity functions (Daniely and Malach, 2020; Cornacchia and Mossel, 2023). The persistent problem of noisy labels, a known challenge in self-play RL, further exacerbates the difficulty for the PV-NN to learn an accurate parity function (Zhou and Riis, 2023). Collectively, these complexities represent a fundamental and significant challenge to contemporary reinforcement learning algorithms in their pursuit of mastering Nim. Our results empirically demonstrate that in a real-world self-play RL setting, the parity function presents a more substantial obstacle for an RL agent to achieve expert-level proficiency in Nim. Given that all impartial games can be theoretically translated into an equivalent Nim game, these identified challenges naturally extend to the broader class of impartial games.
5.3 Controlled experiments
The findings in Section 5.2 raise the question of whether the challenges of AlphaZero-style RL on impartial Nim are fundamentally caused by limitations of reinforcement learning (e.g., insufficient exploration, optimisation instability, data sample out-of-distribution during training), or instead by the difficulty neural networks face in learning the parity-based winning strategy required for optimal play. To disentangle these factors, we design two controlled experiments on small game boards where the entire state space is tractable and repeatedly revisited, thereby removing exploration difficulty, function underfitting due to limited data, and instability in value backpropagation as confounding variables.
The first experiment investigates whether the difficulty lies specifically in neural network function approximation. We replace the neural policy and value networks with tabular lookup tables that can represent the optimal mapping exactly without any inductive bias assumptions. If the tabular tables learn successfully while the neural networks struggle, this would isolate the challenge to neural network generalisation, rather than RL dynamics or game complexity.
The second experiment examines whether the difficulty is intrinsic to the Nim strategy itself. We construct a structurally similar partisan variant of Nim whose optimal strategy does not rely on learning a parity function. If AlphaZero with PV-NNs succeeds in this modified game under identical training conditions, this would further indicate that the challenge in classical Nim originates from the necessity of learning a non-associative, non-linear arithmetic function, not from our modelling choices, optimisation procedure, hyperparameters, or representation.
5.3.1 Tabular AlphaZero-style learning on impartial Nim
To investigate whether the generalisation difficulties of neural networks in learning parity contribute to the challenges faced by AlphaZero-style reinforcement learning on Nim, we implemented a tabular variant of the algorithm where we replaced the PV-NNs with two tabular lookup tables representing the policy and value functions, respectively. This tabular variant serves as a baseline that circumvents neural network approximation errors, allowing us to isolate the impact of function approximation on learning and generalisation in the parity-based Nim environment.
For each state $s$ in the state space, the policy table is initialised to assign equal probability to all legal actions, while the value table is initialised to zero. Given a target policy $\bm{\pi}_{\text{target}}(\mathbf{a}|s)$ for a state $s$ obtained from MCTS, the policy table is updated using a moving average:
$\bm{\pi}(\mathbf{a}|s)\leftarrow(1-\eta)\,\bm{\pi}(\mathbf{a}|s)+\eta\,\bm{\pi}_{\text{target}}(\mathbf{a}|s),$ (13)
where $\eta\in(0,1)$ is a learning rate controlling the update magnitude. After each update, the policy is normalised to ensure it forms a valid probability distribution. Similarly, given a target value $v_{\text{target}}(s)$ from MCTS, the value table is updated via
$v(s)\leftarrow\text{clip}\big((1-\eta)\,v(s)+\eta\,v_{\text{target}}(s),-1,1\big),$ (14)
where $\text{clip}(\cdot,-1,1)$ ensures the value remains within range as used in AlphaZero.
Figure 16: Prediction accuracy of the tabular policy and value models versus the number of distinct states explored during the Tabular AlphaZero-style RL training on the five-heap impartial Nim.
We applied the Tabular AlphaZero-style algorithm to the five-heap Nim game while keeping all algorithmic and training configurations identical to those used in Section 5.2, ensuring a strictly controlled comparison. Figure 16 reports the learning progress, measured by the policy and value prediction accuracy as a function of the number of distinct states visited across 500 training iterations. Both the tabular policy and value estimates improve steadily as more states are encountered, eventually achieving 8080% accuracy on the states in the state space. Notably, the tabular policy accuracy is approximately twice that achieved by the policy network (see Fig. 8(a)). This demonstrates that, on this small Nim board, the challenge does not stem from the AlphaZero training framework itself, but rather from the difficulties of neural networks in learning parities.
Interestingly, the gap between the tabular lookup table and the value network is substantially smaller for value estimation. The final value accuracy of the tabular value closely matches that of the corresponding value network (Fig. 10). This suggests that value prediction, a scalar regression task, is considerably easier to generalise across unseen states than policy prediction, which requires learning a high-dimensional, parity-structured mapping over combinatorial action spaces. Taken together, these results indicate that, within this small five-heap Nim setting, the problem is not due to insufficient exploration, but arises from the difficulty neural networks have in representing the parity-based structure required for accurate policy generalisation.
5.3.2 AlphaZero-style learning on Partisan Nim
To examine whether the learning difficulty observed in mastering Nim primarily arises from the parity issue, we introduce a partisan variant of the game in which each player controls a disjoint subset of heaps and is only allowed to remove matches from those heaps. A player loses if no matches are available on the set of heaps they can remove matches from. Unlike classical Nim, this partisan setting eliminates the need for parity reasoning. Each player’s winning potential depends solely on the total number of matches available within their own designated heaps. As a result, the optimal policy becomes strictly additive and memoryless. Both players should remove exactly one match per turn, thereby maximising the number of future turns they can take. Under optimal play, the winner is the player with the larger initial sum of matches.
We evaluated this setting using the same initial board configuration on five heaps [1,3,5,7,9]\,[1,3,5,7,9]\, as in our impartial Nim experiments. The only modification lies in the move constraints: Player 1 can remove matches exclusively from heaps labelled $a$, $c$, and $e$, while Player 2 is restricted to heaps $b$ and $d$. All algorithmic and representation settings are kept identical to the baseline experiments, allowing us to isolate the effect of removing parity-based reasoning from the learning problem.
Within ten self-play training iterations, the policy network rapidly converges to the theoretically optimal strategy. The trained policy network assigns its highest action probability to removing exactly one match from any legal heap, regardless of which player is to move. For example, in a mid-game position [1,3,5,5,2]\,[1,3,5,5,2]\, where both players have the same number of matches, the best move for each player is to remove one match from one of the designated matches. When this position is labelled as Player 1’s turn, the policy network assigns probabilities of $0.53$, $0.29$, and $0.11$ to removing one match from heaps $e$, $c$, and $a$, respectively. When labelled as Player 2’s turn, it assigns $0.64$ and $0.19$ to actions $b1$ and $d1$, respectively, again matching the optimal strategy.
This stark contrast between the agent’s performance on impartial Nim (which struggled significantly) and its quick mastery of partisan Nim supports our central finding that the inherent difficulty for generic neural networks to implicitly model abstract, non-associative functions like parity is the primary bottleneck in this learning paradigm. Furthermore, its exceptional performance using the same board representation as in the impartial Nim experiments suggests that the game state representation itself does not adversely impact performance when the underlying optimal function is simpler and more associative.
5.4 Limitations and future work
To contextualise this, consider our analysis of a specific Chess position (Fig. 5) which subtly mirrors a rudimentary parity (nim-like) configuration. Our investigation revealed that even the highly optimised and extensively trained LCZero, which for all practical purposes embodies the AlphaZero paradigm, struggled to accurately assess this particular position. This observation provides a compelling precursor to our central argument: in Nim, where the winning strategy is directly and inextricably linked to the parity function (nim-sum), this issue becomes dramatically more pronounced. The core difficulty lies in the generic neural network’s struggle to implicitly learn a perfect parity function from noisy, self-play generated data, especially as the input dimensionality increases. This inherent representational bottleneck for combinatorial logic, distinct from the complexity of typical board game patterns, is what we aim to thoroughly explore in this work.
Our current approach employs an LSTM network over a structured representation of Nim states. While this provides a simple and flexible sequence-learning interface that can accommodate inputs of arbitrary length, which is consistent with the nature of the parity function, it does not explicitly capture the combinatorial structure or the permutation invariance inherent to impartial games. In Nim, permuting heap order does not alter optimal play, yet our representation and LSTM implicitly introduce order sensitivity, yielding a potential inductive bias mismatch between architecture and game semantics. More structure-aware neural networks, such as permutation-invariant encoders (e.g., Deep Sets (Zaheer et al., 2017)), or Transformers without positional encodings (Haviv et al., 2022) may be better aligned with the symmetry and parity properties of impartial games, and could improve sample efficiency and generalisation. As we have surveyed in Section 2.2, these established neural architectures still face intrinsic challenges in modelling a perfect parity function, especially from long binary strings. We are therefore interested in evaluating alternative neural architectures as PV-NNs for learning impartial games, and we leave this to future work.
Although we adopt the standard AlphaZero framework, our study does not evaluate several notable extensions that could potentially strengthen performance. For example, Gumbel AlphaZero (Danihelka et al., 2022) has demonstrated improved exploration and policy improvement guarantees, and recent AlphaZero-derived solving frameworks (Wu et al., 2022, 2023; Dantsin et al., 2023) explicitly target exact reasoning over full game graphs. Exploring such variants, particularly in settings where theoretical optimal play is known and parity-based reasoning is central, represents an important direction for future investigation. A more comprehensive empirical comparison across these methods may yield deeper insight into the challenges identified in this work. We leave this for future work.